OpenStudy (anonymous):

find the equation of the line that contains the point whose coordinates are (2,1) and has slope -2

OpenStudy (anonymous):

IDK if you have learned point slope form... but you can use that form (y-y1) = m(x-x1) and solve for y. If you have to do it the other way, plug in your x and y coordinates in for Y = mX +b. Plug in -2 for your slope and then solve for your b. After that just write y = (your slope)x+(the b value you get).